OpenStudy (anonymous):

write an equation for the line that passes through the point with the given slope: (-3,-4), m =3

OpenStudy (anonymous):

(y+4)/(x+3) = 3

OpenStudy (anonymous):

y=3x+5

OpenStudy (anonymous):

Both are correct; him1618 did slope = y-y1/(x-x1) = 3 and adanne1990 did the algebra to arrive at the answer

OpenStudy (anonymous):

y=3x+5

OpenStudy (anonymous):

how can you show the work?

OpenStudy (anonymous):

i need a white board

OpenStudy (anonymous):

wait ...

OpenStudy (anonymous):

k